

********************************************************************
UARS WEEKLY STATUS REPORT - 17 JUN 2005
********************************************************************
***********
SOLSTICE
***********
SOLSTICE ended day-only operations on day 159. The instrument is in
good health gathering solar data on all available orbital tracking
periods. On average, we received 64% of each day's data amounting to
about 719 minutes of solar observing time per day.
***********
SUSIM
***********
During this past week, SUSIM has successfully gathered its daily mid
(1.1 nm) resolution, MgII, and continuous monitoring scans for time
periods with UARS telemetry or tape recording. On June 15, a full
0.15 nm (high) resolution scan was also performed.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005049 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005039 has been processed and is waiting for DAAC archival. MEPS data for
2005040 to 2005049 is currently being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 04 June 2005 (Orbit 75151, GMT Day 155) through
10 June 2005 (Orbit 75256, GMT Day 161).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 155/1305 75159 POWER ON
SUSIM 157/1538 75191 POWER ON
SOLSTICE 159/1348 75220 DAY ONLY TO FULLTIME
The beta angle ranged between the angles of 20.9 degrees to 46.3
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 19 June 2005 (DOY 170).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 6.13 to 5.43 degrees C. Additional
battery performance data for each battery on 03 June 2005 and 10 June
2005 are provided in Appendix D.
The clock error ranged from 09.9 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 02
SUSIM = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to closely monitor the status of the
batteries. The batteries maintained their VT-6 charge level
throughout the week, until 09 June 2005 on the 1630 event. At that
time the VT level was dropped down to VT-5 in order to stop the
increasing battery 2 temperature. The batteries were able to support
HALOE, SUSIM and SOLSTICE fulltime at this lower charge rate and the
temperature came down 2.7 degrees in less than a day.
The UARS OBC clock rollover effort continued this week with the
completion of procedure. Testing of this procedure was performed on
the UARS simulator with very good preliminary results. There were
only two minor problems encountered. Next week the procedure will be
updated to correct these problems and additional telemetry validation
will be added. The uplink to the spacecraft is scheduled for 02
August 2005.
MEETINGS
-None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 July 2005 (DOY 191). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 June 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
75159 155/1305 05-180 HALOE POWER ON - PART ONE
75162 155/1803 05-181 HALOE POWER ON - PART TWO
75168 156/0244 05-182 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
75191 157/1538 05-183 SUSIM POWER ON - PART ONE
75220 159/1348 05-184 SOLSTICE DAY ONLY TO FULLTIME OPS
75228 160/0230 05-185 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
75235 160/1316 05-186 MPS VT LEVEL 5
75246 161/0659 05-187 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 June 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 June 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
75168 156/0245 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
75228 160/0234 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
75246 161/0702 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 June 2005
03 June 2005 - GMT Day 154 Beta = 18.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 11.2/ -22.4 -3.68/-4.94 +00.0/+00.0 05.8 0.00 0.0
2 151.2/ -89.6 +8.73/+7.74 +35.2/-18.4 26.1 1.04 23.4
3 -22.4/-700.0 +2.60/+1.65 +17.6/-09.6 26.1 1.08 10.5
10 June 2005 - GMT Day 161 Beta = 46.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 00.0/ -22.4 -3.68/-4.94 +00.0/+00.0 05.8 0.00 0.0
2 -05.6/-179.2 +8.73/+7.09 +30.4/-21.6 25.4 1.04 23.8
3 00.0/-700.0 +2.60/+1.66 +16.8/-11.6 25.4 1.05 12.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
75237 160/1316 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 10 June 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-030 FOT 75191 157/1538 30832 Generic Late Acquisition
05-030 FOT 75228 160/0328 30832 Generic Late Acquisition
05-030 FOT 75244 161/0439 30832 Generic Late Acquisition
05-030 FOT 75245 161/0549 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-031 FOT 75122 153/0111 39664 Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-030 FOT See List June 05 30832 Generic Late Acquisition